You observe the latest dividend paid of $4 per share. The growth rate is projected to...

  1. You observe the latest dividend paid of $4 per share. The growth rate is projected to be a constant 5% per year. Your required rate of return is 8%.

    a. What price are you willing to pay for that stock today?

    b. What is expected stock price in 1 year?

    c. Find dividend yield.

    d. Find capital gains.
Answer #1

Pirce today = D1=(Ke- g) $140.00 (4*1.05)/(0.08-0.05) Price after 1 year = $147.00 (140*1.05] Dividend yield = dividend price